OpenStudy (tootzrll):

Two cars traveled equal distances in different amounts of time. Car A traveled the distance in 2 h, and Car B traveled the distance in 1.5 h. Car B traveled 15 mph faster than Car A. How fast did Car B travel?

OpenStudy (bluecreep):

I need more info than this to answer but right now it would seem that car A is going half the speed because if you were car A and you were going to travel 80 miles and you wanted to get there in a our you would need to go 80 mph but if you were needed to i 2 hours you would have had to go 40 mph and since car b made it in 1.5 hours it had to have gone 3 quarters of the speed.

OpenStudy (tootzrll):

thanx :]]

OpenStudy (matt101):

You can solve this with the information given! Just form some equations with algebra! Let's say the speed of Car A is v. That means the speed of Car B is v+15 (it traveled 15 mph faster than Car A). We also know that both cars traveled equal distances, d, and that t=2 for Car A and t=1.5 for Car B. Using v=d/t, we can come up with equations to describe both cars: \[Car \space A: v=\frac{d}{2}\]\[Car \space B: v+15=\frac{d}{1.5}\] Two equations, two unknowns. We're interested in solving for a speed, so you just need to rearrange both equations to solve for d, and because d is equal in both cases, you can set those rearranged equations equal to one another and solve for v! Keep in mind though that we defined the speed of Car B as v+15!